OpenStudy (cutiekawdfi):

help pls.. 1. identify the vertex, axis of symmetry, max/min, and domain and range. f(x)= -(x-3)^2 + 2 2. write in vertex form. f(x)=8x^2+8x-12

OpenStudy (jango_in_dtown):

hi @cutiekawdfi

OpenStudy (cutiekawdfi):

hi

OpenStudy (jango_in_dtown):

when you have y=a(x-b)^2+c, the vertex is (b,c). now try to put the first equation in that form.

OpenStudy (jango_in_dtown):

y=f(x)=-(x-3)^2+2=(-1){x-3}^2+2 so here a=-1,b=3,c=2 the vertex is (b,c)=(3,2)

OpenStudy (jango_in_dtown):

axis of symmetry is x=3

OpenStudy (jango_in_dtown):

max value=2

OpenStudy (jango_in_dtown):

domain all real numbers

OpenStudy (cutiekawdfi):

what about the second one

OpenStudy (cutiekawdfi):

@jango_IN_DTOWN